Kabbed:

Your boss hands you the monthly data that show the number of orders coming in to and out of the warehouse. The data are in the table below. Explain to your boss, in complete sentences, the solution to this system and what the solution represents.

Kabbed:

http://prntscr.com/n9bqy4

Kabbed:

"Oi! I have a few minutes before I need to leave but I'll try to show you how to answer your question: Your month is your x-coordinate The no. of orders IN is your y-coordinate. For example (1, 3) Likewise, we will have another set of points where Your month is your x-coordinate The no. of orders OUT is your y-coordinate You will first need to calculate the lines of each set of points. Start by first finding the slope. \(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\) To find the slope, you only need 2 points. After you find the slope, plug that into y = mx + b where m is the slope and b is the y-intercept So you should have a number replacing 'm' After that, plug in one of your points (it can be one of the points you used to calculate your slope) and plug that into the equation and solve for b Once you get your two equations for both data the orders OUT and the orders IN, you set them equal to each since they both will be in the form of y=mx+b And so you can solve for x which would be the solution to the system" This is what was said to me by tso but sadly I still don't understand

Kabbed:

Slope = I don't know what to do or what numbers to write

Kabbed:

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\)

Kabbed:

@Hero

Hero:

You are given two sets of points to plot: The first set of points is represented by \((m, i)\) where \(m\) represents the month and \(i\) represents orders coming in. The first set of points are (1, 3), (2, 6), (3, 9) and (4,12) The second set of points is represented by \((m, o)\) where \(m\) represents the month and \(o\) represents orders going out. The second set of points are (1,3), (2,4), (3,5), (4, 6) Plot each set of points on graph paper. Draw a line through the first set of points. Then draw a line through the second set of points. The intersection point of the two lines is the solution to the system.

Hero:

@Kabbed are you alright?

Kabbed:

So sorry I left but I've returned :) Thank you so much for helping me

Kabbed:

so the answer is (1,3) because that's what I found out after graphing thats where the two lines met

Kabbed:

@hero

Hero:

Correct

Kabbed:

Hey thank you I messaged you what I wrote, is it good lol

Hero:

The only thing that matters is that you understand how to do the problem on your own.

Hero:

I hope I was able to help you understand.

Kabbed:

so the solution represents the break even point not profit

Kabbed:

does everyone say break even point for these types of problems

Hero:

It represents when the orders in = orders out

Hero:

But of course that occurs in the month of January.

Kabbed:

Oh okay I see

Hero:

Basically, when you look at it properly, it just means they shipped out and received three orders in the first month.

Hero:

That's how to properly sum up the meeting point of both lines.

Kabbed:

thank you